Circular No. 7885

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 S5 T. Mäkinen, Finnish Meteorological Institute, Helsinki, reports the

discovery of another apparent comet (cf. IAUC 7327) from the second systematic survey of hydrogen Lyα emission appearing on SWAN images (wavelength range 10–180 nm) taken from the SOHO spacecraft during September-November 2000 (discovery observation below). Mäkinen notes that all comets brighter than total visual mag ≈ 11 can be seen on the SWAN full-sky maps. The stated uncertainty in the reported positions is 0^{o}.5 due to the resolution of the SWAN instrument.

2000 UT  α_{2000}  δ_{2000}  m_{1}
Sept. 19.658  11^{h}35.^{m}9  −12^{o}13′  11.6

Further positional and orbital information, together with an ephemeris to encourage searches of archival optical images, are given on MPEC 2002-H41. Derived total visual magnitudes from the Lyα fluxes (using a formula by Jorda et al. 1992, Asteroids, Comets, Meteors 1991, 285): 2000 Sept. 21.77 UT, 10.8; Oct. 1.42, 10.0; 5.23, 9.4; 7.35, 8.4; 10.38, 7.7; 12.45, 7.3; 14.37, 7.5; 17.40, 7.8; 21.38, 7.9; Nov. 4.42, 8.2.

SUPERNOVAE 2002cl, 2002cm, 2002cn, 2002co G. Altavilla and S. Benetti, Osservatorio di Padova, and their col-

leagues (cf. IAUC 7780) report the discovery of four supernovae on CCD frames taken on Apr. 7/8 (via comparison with reference images having limiting mag ∼ 24 taken with the same equipment at previous observing runs) during their search with the ESO/MPI 2.2-m telescope.

SN  α_{2000}  δ_{2000}  V  Offset  Type z
2002cl 13^{h}44^{m}09.^{s}94 − 0^{o}12′57′′.8 21.6 2′′.7 W, 1′′.4 S Ic 0.072
2002cm 13 52 03.67 −11 43 08.5 22.3 1′′.6 W, 6′′.1 N II 0.087
2002cn 13 53 21.24 −12 15 29.5 22.3 0′′.8 W, 0′′.7 N Ia 0.302
2002co 14 10 53.04 −11 45 25.0 22.6  —  II 0.318

No parent galaxy was visible in the reference image of SN 2002co. Spectra obtained on Apr. 21 with the ESO 8.2-m KUEYEN Very Large Telescope (+ FORS2) reveal the types and redshifts listed above. From the spectral features, it is estimated that SN 2002cl was ∼ 3 weeks after maximum, SN 2002cm was several weeks after maximum, and SN 2002cn was a few days after maximum. SN 2002co shows Balmer emissions similar to those of SN 1996L when close to maximum (Benetti et al. 1999, MNRAS 305, 811).

2002 April 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green